OpenStudy (anonymous):

I need this problem broke down to it's simplest form. 9357/2520

OpenStudy (paxpolaris):

3 goes into both

OpenStudy (paxpolaris):

9+3+5+7=24..............2+5+2+0=9

OpenStudy (anonymous):

1/4x2/5x3/6x4/7x5/8x9/10 please work this problem for me

OpenStudy (paxpolaris):

so, \[{9357\over2520} = {3119 \over 840}\] although this may not be the simplest form... for that factor 840, because it's easier...\[840 = 84 \times 10 = 12 \times 7 \times 10\]\[ = \Large 2^3\times 3 \times 5 \times 7\] 2, 3, and 5 don't go into 3119, so try 7... it doesn't go either... so final anwer is 3119/840

OpenStudy (paxpolaris):

\[\Large\frac14\times \frac25 \times \frac36 \times \frac47\times \frac58\times \frac9{10}\]

OpenStudy (paxpolaris):

\[=\frac14\times \frac25 \times \frac {\cancel3} {\cancel62} \times \frac47\times \frac58\times \frac9{10}\]

OpenStudy (paxpolaris):

\[=\frac14\times \frac{\cancel2}5 \times \frac1{\cancel2} \times \frac47\times \frac58\times \frac9{10}\]

OpenStudy (paxpolaris):

\[\frac1{\cancel4}\times \frac1{\cancel5} \times 1 \times \frac{\cancel4}7\times \frac{\cancel5}8\times \frac9{10}\]

OpenStudy (paxpolaris):

\[= \Large {9 \over 7\times 8 \times 10}\]